# Preconditioning without a preconditioner: faster ridge-regression and Gaussian sampling with randomized block Krylov subspace methods

### v1

@@ 6 Gaussian sampling
Figure 4: Maximum sample error versus matrix-loads for Lanczos square root (Refer to caption) and block-Lanczos square root (Refer to caption).
Several applications in data science and statistics require sampling Gaussians with a given mean and covariance (Thompson, 1933; Besag et al., 1991; Bardsley, 2012; Wang & Jegelka, 2017). A standard approach is to transform an isotropic Gaussian vector. Indeed, suppose $\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (i.e. that entries of $\mathbf{b}$ are independent standard Gaussians). Then, for positive definite $\mathbf{A}$,
$\bm{\mu}+\mathbf{A}^{1/2}\mathbf{b}\sim\mathcal{N}(\bm{\mu},\mathbf{A}).$ (24)
@@
Recall the expression Equation 4:
$\mathbf{A}^{1/2}\mathbf{b}=\frac{2}{\pi}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}\,\mathrm{d}z.$ Equation 4
@@
This and closely related methods appear throughout the literature (Chow & Saad, 2014; Pleiss et al., 2020; Chen, 2024).
@@
Both algorithms can be efficiently implemented using the (block)-Lanczos algorithms; see Section C.3.
Using our analysis of augmented block-CG we are able to prove the following convergence guarantee.
@@
$C=\log(16\lambda_{1}/\lambda_{d})$, and
$\varepsilon_{t}:=C\exp\left(-\frac{t-(3+\log(d))}{3\sqrt{(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)}}\right)$
@@
$\Bigg{\{}\forall i:\frac{\|\mathbf{A}^{1/2}\mathbf{b}^{(i)}-\mathsf{bsq}_{t}^{(i)}\|}{\|\mathbf{A}^{1/2}\|\|\mathbf{b}^{(i)}\|}\leq\varepsilon_{t}\Bigg{\}}.$
The proof is contained in Appendix C and basically amounts to integrating Corollary 3.4. In particular, when sampling $m$ Gaussians, $\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(m)}$ are independent isotropic Gaussian vectors. Therefore, $[\mathbf{b}^{(2)},\ldots,\mathbf{b}^{(m)}]$ is a Gaussian matrix independent of $\mathbf{b}^{(1)}$, and hence $\mathsf{bcg}_{t}^{(1)}(z^{2})$ will enjoy the benefits of augmented block-CG described in this paper. Thus, we expect $\mathsf{bsq}_{t}^{(1)}(z^{2})$ to converges more quickly the larger $m$ is. Critically, since $\mathsf{bsq}_{t}^{(i)}(z^{2})$ are identically distributed, this acceleration is also enjoyed by $\mathsf{bsq}_{t}^{(i)}(z^{2})$ for all $i$, without the need to generate any new random vectors!
Thus, block-Lanczos square root method allows us to sample $m$ Gaussians with roughly
$m\Big{(}\log(d)+\sqrt{\lambda_{r+1}/\lambda_{d}}\log\big{(}\log(\lambda_{1}/\lambda_{d})/\varepsilon\big{)}\Big{)}$ (25)
@@
matrix-vector products, where $r=O(m/\log(m))$. In contrast, existing bounds for methods like the single-vector Lanczos square root method Definition 6.1 (Chen et al., 2022; Pleiss et al., 2020) sample a single Gaussian vector with roughly $\sqrt{\lambda_{1}/\lambda_{d}}\log(1/\varepsilon)$ matrix-vector products. Therefore, when $\lambda_{r+1}\ll\lambda_{1}$, the total number of matrix-vector products is reduced significantly by using the block method in Definition 6.2.
6.1 Numerical experiments
@@
We perform an experiment comparing the block-Lanczos square root iterate to the standard Lanczos square root iterate. In particular, in Figure 4, we compare the maximum error in approximating $\mathbf{A}^{1/2}\mathbf{b}^{(i)}$, $i=1,\ldots,m$ for the two methods. All runs of the single-vector method are done in parallel so that the number of matrix-loads is independent of $m$. The block method outperforms the single-vector method, and the performance gains are often significant.
@@ 7 Outlook
We have introduced a variant of the block-CG method which outperforms Nyström PCG, both in theory and in practice. Our work is perhaps the first to provide theoretical evidence of the virtues of block Krylov subspace methods for solving a single linear system of equations and sampling Gaussian vectors. There is significant potential to derive better theory and algorithms for such methods, particularly by developing a better understanding of randomized low-rank approximation (Meyer et al., 2024).
In order to derive our theoretical bounds we describe new variants of Nyström PCG that use a block-Krylov Nyström approximation and prove theoretical guarantees for these variants. Our theory and numerical experiments suggest that these methods perform better than the simple Nyström approximation suggested in (Frangella et al., 2023).
Finally, it would be interesting to understand the extend to which similar ideas can be extended to other tasks involving matrix-functions, such as estimating the trace of matrix-functions. Existing algorithms often involve applying matrix-functions to many Gaussian vectors (Ubaru et al., 2017; Epperly et al., 2024; Meyer et al., 2024), so it seems likely that using block-Krylov methods may be advantageous to using single-vector methods. However, existing theory does not reflect this.
@@ Acknowledgements
We thank Ethan Epperly for useful discussions and feedback.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v2

@@ 6 Gaussian sampling
Several applications in data science and statistics require sampling Gaussians with a given mean and covariance [59, 3, 2, 63]. A standard approach is to transform an isotropic Gaussian vector. Indeed, suppose $\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (i.e. that entries of $\mathbf{b}$ are independent standard Gaussians). Then, for positive definite $\mathbf{A}$,
(32) $\bm{\mu}+\mathbf{A}^{1/2}\mathbf{b}\sim\mathcal{N}(\bm{\mu},\mathbf{A}).$
@@
Recall the relation:
Eq. 4 $\mathbf{A}^{1/2}\mathbf{b}=\frac{2}{\pi}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}\,\mathrm{d}z.$
@@
This and closely related methods appear throughout the literature [12, 52, 7].
@@
Both algorithms can be efficiently implemented using the (block)-Lanczos algorithms; see e.g. [7].
We can use the bounds for augmented block-CG from Section 4 to derive bounds for the block-Lanczos square root algorithm. In particular, using Theorem 4.2, we prove:
@@
$\Bigg{\{}\forall i:\frac{\|\mathbf{A}^{1/2}\mathbf{b}^{(i)}-\mathsf{bsq}_{t}^{(i)}\|}{\|\mathbf{A}^{1/2}\|\|\mathbf{b}^{(i)}\|}\leq\log\left(16\kappa\right)\exp\left(-\frac{t-(2+\log(d)/2)}{3\sqrt{\kappa_{r+1}(0)}}\right)\Bigg{\}}.$
Similarly, using Theorem 4.3, we prove:
Theorem 6.4.
Let $\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(m)}\in\mathbb{R}^{d}$ be independent standard Gaussian random vectors and $r\geq 0$. Then, for some
$s=O\left(\frac{r}{m}\log(\Delta)+\log\left(\frac{d}{\delta}\right)\right),\quad\Delta:=\min_{i=1,\ldots,(m-1)\lceil r/(m-1)\rceil-1}\frac{\lambda_{i}-\lambda_{i+1}}{\lambda_{1}},$
with probability at least $1-\delta$,
$\Bigg{\{}\forall i:\frac{\|\mathbf{A}^{1/2}\mathbf{b}^{(i)}-\mathsf{bsq}_{t}^{(i)}\|}{\|\mathbf{A}^{1/2}\|\|\mathbf{b}^{(i)}\|}\leq\log\left(16\kappa\right)\exp\left(-\frac{t-s}{3\sqrt{\kappa_{r+1}(0)}}\right)\Bigg{\}}.$
Proof 6.5 (Proof of Theorems 6.3 and 6.4).
Our proof is optimized for simplicity rather than sharpness. For notational convenience, let $\mathbf{b}=\mathbf{b}^{(1)}$ and $\mathbf{\Omega}=[\mathbf{b}^{(2)},\ldots,\mathbf{b}^{(m)}]$. As noted in Section 6, $\mathbf{\Omega}\in\mathbb{R}^{d\times(m-1)}$ is independent of $\mathbf{b}$. Consider the event
(33) $\Bigg{\{}\forall\mu\geq 0:\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq\varepsilon_{t}^{\prime}(\mu)\Bigg{\}},$
where
(34) $\varepsilon_{t}^{\prime}(\mu)=2\exp\left(-\frac{t-s}{3\sqrt{\kappa_{r+1}(\mu)}}\right).$
To prove Theorem 6.3, note that the choice of $r$ ensures that $m-1\geq(r+2)q$ for $q=\lceil\log(m/\delta)/\log(100)\rceil$. Therefore, with $s=2+\log(d)/2$, Theorem 4.2 guarantees that Eq. 33 holds with probability at least $1-\delta/m$. Likewise, to prove Theorem 6.4, note that the Theorem 4.3 guarantees that Eq. 33 holds with probability at least $1-\delta/m$ (where we have used that $m\leq d$ so that $\log(dm/\delta)=O(\log(d/\delta))$).
From this point on we condition on Eq. 33, and the two proofs are identitcal. Applying standard norm inequalities we obtain a bound
(35) $\|\mathbf{A}\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(\mu)\|$ $=\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\mathbf{A}_{\mu}^{1/2}(\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu))\|$
(36) $\leq\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\|\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}.$
Under the assumption the event in Eq. 33 holds, we bound
(37) $\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}\leq\varepsilon_{t}(\mu)\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}\leq\varepsilon_{t}(\mu)\|\mathbf{A}_{\mu}^{-1/2}\|\|\mathbf{b}\|.$
We observe that
(38) $\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\|=\max_{i}\frac{\lambda_{i}}{\sqrt{\lambda_{i}+\mu}}=\frac{\lambda_{1}}{\sqrt{\lambda_{1}+\mu}}$
and similarly
(39) $\|\mathbf{A}_{\mu}^{-1/2}\|=\max_{i}\frac{1}{\sqrt{\lambda_{i}+\mu}}=\frac{1}{\sqrt{\lambda_{d}+\mu}}.$
(40) $\|\mathbf{A}\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(\mu)\|\leq\frac{\lambda_{1}\varepsilon_{t}^{\prime}(0)\|\mathbf{b}\|}{\sqrt{(\lambda_{1}+\mu)(\lambda_{d}+\mu)}}.$
Therefore, applying the triangle inequality for integrals and Eq. 40 we compute a bound
(41) $\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{bsq}_{t}^{(1)}\|$ $=\frac{2}{\pi}\biggl{\|}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(z^{2})\mathrm{d}{z}\biggr{\|}$
(42) $\leq\frac{2}{\pi}\int_{0}^{\infty}\|\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(z^{2})\|\mathrm{d}{z}$
(43) $\leq\frac{2}{\pi}\int_{0}^{\infty}\frac{\lambda_{1}\varepsilon_{t}^{\prime}(0)\|\mathbf{b}\|}{\sqrt{(\lambda_{1}+z^{2})(\lambda_{d}+z^{2})}}\mathrm{d}{z}.$
A direct computation reveals
(44) $\int_{0}^{\infty}\frac{1}{\sqrt{(\lambda_{1}+z^{2})(\lambda_{d}+z^{2})}}\mathrm{d}{z}=\frac{K(1-\lambda_{1}/\lambda_{d})}{\sqrt{\lambda_{d}}},$
(45) $\frac{2}{\pi}\sqrt{x}K(1-x)\leq\frac{5}{4\pi}\log(16x)\leq\frac{1}{2}\log(16x).$
Therefore, applying Eq. 45 with $x=\kappa=\lambda_{1}/\lambda_{d}$ to Eqs. 43 and 44 we get a bound
(46) $\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{bsq}_{t}^{(1)}\|\leq\frac{\sqrt{\lambda_{1}}}{2}\log\left(16\kappa\right)\varepsilon_{t}^{\prime}(0)\|\mathbf{b}\|.$
To obtain the final result, we observe that $\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(m)}$ and $\mathsf{bsq}_{t}^{(1)},\ldots,\mathsf{bsq}_{t}^{(m)}$ are permutation invariant in distribution. Thus, the above result in fact applies to each $\mathbf{b}^{(i)}$ and $\mathsf{bsq}_{t}^{(i)}$ pair with probability at least $1-\delta/m$. Applying a union bound over these $m$ events gives the main result.
Theorems 6.3 and 6.4 gives bounds on the matrix-vector products required by the block-Lanczos method. For instance, Theorem 6.3 implies we can use roughly
(47) $m\Big{(}\log(d)+\sqrt{\kappa_{r+1}}\log\big{(}\log(\kappa)/\varepsilon\big{)}\Big{)}$
@@
matrix-vector products, where $r=O(m/\log(m))$. In contrast, existing bounds for methods like the single-vector Lanczos square root method Definition 6.1 [9, 52] sample a single Gaussian vector with roughly $\sqrt{\kappa}\log(1/\varepsilon)$ matrix-vector products. Therefore, when $\lambda_{r+1}\ll\lambda_{1}$, the total number of matrix-vector products is reduced significantly by using the block method in Definition 6.2.
Remark 6.6.
For sampling Gaussians, some preconditioning-like type approaches can be applied used [12, 52, 26]. As far as we are aware, there are no theoretical guarantees for such methods similar to the ones presented in the current work. In addition, analogous to the case of linear systems, we expect block methods to enjoy the benefits of working over a larger subspace.
Remark 6.7.
Relating KSMs for matrix functions to KSMs for (shifted) linear systems via integral relations widely used in order to design and analyze algorithms for matrix functions; see e.g. [25, 57, 9, 64]. However, it is generally difficult to use preconditioners for linear systems for matrix functions, as preconditioned Krylov subspaces do not generally satisfy a shift invariance property. As such, the integral relation no longer yields an efficient to implement algorithm. On the other hand, we expect similar analyses, based on implicit preconditioning with block-CG, to work for other functions.
6.1 Sampling with the inverse covariance
Computing $\mathbf{A}^{-1/2}\mathbf{b}$ is used to move to a “whitened” coordinate space, and has found use in a number of data-science applications [37, 52].
Definition 6.8.
Note that $\mathsf{bisq}_{t}^{(i)}=\mathbf{A}^{-1}\mathsf{bsq}_{t}^{(i)}$. Therefore, Theorem 6.3 immediately gives a bound for Definition 6.8 since
(48) $\|\mathbf{A}^{-1/2}\mathbf{B}-\mathsf{bisq}_{t}\|$ $=\|\mathbf{A}^{-1/2}\mathbf{B}-\mathbf{A}^{-1}\mathsf{bsq}_{t}\|$
(49) $=\|\mathbf{A}^{-1}(\mathbf{A}^{1/2}\mathbf{B}-\mathsf{bsq}_{t})\|$
(50) $\leq\|\mathbf{A}^{-1}\|\|\mathbf{A}^{1/2}\mathbf{B}-\mathsf{bsq}_{t}\|.$
7 Numerical experiments
We use the block-Lanczos algorithm with full-reorthogonalization in order to implement block-CG and the block-Lanczos square root iterates. Code to replicate the experiments in this paper is available at https://github.com/tchen-research/precond_without_precond. The $s$ matrix-loads required to build the Nyström preconditioner are accounted for in the plots. For Nyström PCG, we set $\theta$ as the smallest eigenvalue of the Nyström approximation. We use the same random Gaussian matrix $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ for Nyström PCG and block-CG. More numerical experiments, including without full-reorthogonalization, are included in Appendices B and A.
7.1 Convergence
We begin by comparing the convergence of the methods discussed in this paper on several test problems. The results of this experiment are illustrated in Fig. 2, and as expected, our augmented block-CG method outperforms the other methods, often by orders-of-magnitude. We also observe that Nyström PCG benefits from using $s>1$; i.e. from performing more than one passes over $\mathbf{A}$ when building the preconditioner.
As discussed in Section 1.2, throughout this paper we assume that matrix-loads (iterations) are the dominant cost. For reference, we have also show convergence as a function of matrix-vector products in Fig. 8.
Figure 2: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-loads for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on several test problems.
7.1.1 Intro figure
Figure 1 is run on the fastdecay problem ($d=8000$) with $\ell=10$. Nyström preconditioning is run with a grid of parameters. Specifically, we use depths $s\in\{1,5,11\}$ and a range of $\theta$, including the choice $\theta=\lambda_{\ell}(\mathbf{A}\langle\mathbf{\Omega}\rangle)$.
The timings are produced based on a setting where $\mathbf{A}$ cannot be stored in its entirety in fast memory. In particular, we store $\mathbf{A}$ in 8 $1000\times 8000$ chunks and perform (block) matrix-vector products with $\mathbf{A}$ by sequentially loading a single chunk from the disk into random access memory and performing the appropriate part of the products. Similar to [61, §6.1], we observe that the runtime of the algorithm is dominated by the cost of loading these chunks into random access memory.
7.2 Regularization parameter
In Fig. 3 we plot the error after a fixed number of matrix loads as a function of $\mu$. As expected, this plot indicates that our augmented block-CG outperforms the other methods for each value of $\mu$. However, block-CG reaches its maximum attainable accuracy sooner, stagnating.
Perhaps more importantly, our block-CG method can efficiently compute the solution to Eq. 1 for many values of $\mu$, without the need for additional matrix-loads (see Theorem 3.7). This is in contrast to Nyström PCG which requires a new run for each value of $\mu$. For tasks such as ridge regression, this is of note.
Figure 3: Relative error $\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}_{\mu}}/\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}$ after a fixed number of matrix-loads as a function of the regularization parameter $\mu$ for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption).
7.3 Sampling Gaussians
@@
We perform an experiment comparing the block-Lanczos square root iterate to the standard Lanczos square root iterate. In particular, in Fig. 4, we compare the maximum error in approximating $\mathbf{A}^{1/2}\mathbf{b}^{(i)}$, $i=1,\ldots,m$ for the two methods. All runs of the single-vector method are done in parallel so that the number of matrix-loads is independent of $m$. The block method outperforms the single-vector method, and the performance gains are often significant.
Figure 4: Maximum relative sample error $\max_{i}\|\mathbf{A}^{1/2}\mathbf{b}_{i}-\mathsf{alg}\|/\|\mathbf{A}^{1/2}\mathbf{b}_{i}\|$ versus matrix-loads for Lanczos square root (Refer to caption) and block-Lanczos square root (Refer to caption).
@@ 8 Outlook
We have introduced a variant of the block-CG method which outperforms Nyström PCG in certain settings. Our work provides theoretical evidence of the virtues of block Krylov subspace methods for solving a single linear system of equations and sampling Gaussian vectors. We believe future work should study a more practical variant of augmented block-CG, where deflation is used to reduce the block-size as soon as it is recognized that the benefits of a large block size are no longer justified.
It would also be interesting to understand the extent to which similar ideas can be extended to other tasks involving matrix-functions, such as estimating the trace of matrix-functions. Existing algorithms often involve applying matrix-functions to many Gaussian vectors [62, 21, 44], so it seems likely that using block-Krylov methods may be advantageous to using single-vector methods. However, existing theory does not reflect this.
@@ Acknowledgements
We thank Ethan Epperly, Daniel Kressner, and David Persson for useful discussions and feedback.